# Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:

### v1

@@ 3.1 Binding energy
Firstly, we calculate the optimization energy of 9 ASW clusters shown in LABEL:tab:opt_water along with their optimized structures. It is seen that the ASW(1) cluster is, on average, K more stable than the other 8 ASW clusters. We then calculate the BE of OH radical at the -H binding sites of 9 ASW clusters. All the calculated (+) BSSE BE values are plotted in Fig. 1 and tabulated in LABEL:tab:BE_OH.
@@
Instead of taking the average BE over all the binding sites, we consider the average over the highest -H binding sites in each cluster. This averaging would be more appropriate for estimating the BE at a low temperature (interstellar condition). As a result, we obtain average BE values of K and K for the OH radical with and without considering BSSE corrections, respectively. Depending upon the binding sites, the BEs are rather broadly distributed (see LABEL:tab:BE_OH). To verify the distribution of the BSSE-corrected BE values over the highest -H binding sites in each cluster, we organize them in a distribution plot with a bin width of 500 K shown in Fig. 2. The distribution is well reproduced by an overlay of Gaussian Probability distribution function:
@@
where is the mean and is the standard deviation. The fitted parameters for the BE values distribution are the mean () K and standard deviation () K. It implies that the Boltzmann distribution (uniform distribution to the highest BE site at K) holds for each local region, which is further supported by the fitting using the Boltzmann model equation (Fig. 7). On the contrary, the BE distribution considering all the 75 sites (BSSE-corrected and uncorrected) does not mimic a Boltzmann distribution (non-Gaussian type), as depicted in Fig. 8.
@@
To minimize the computational burden, we perform the calculation of the BE for the remaining 6 radicals (i.e., CH, NH, SH, CN, NS, and NO) by considering only the ASW(1) cluster. This particular ASW cluster is the most stable (lowest energy) among the 9 ASW clusters we have mentioned earlier (see LABEL:tab:opt_water). Then, we select the highest physisorbed BE values among 7 binding sites available in ASW(1) (see Table 7). It is interesting to note that except for NO, all the other 6 radicals (i.e., CH, NH, OH, SH, CN, and NS) give the highest BEs (considering both physisorption and chemisorption) with the fifth -H position of ASW(1) cluster. Since we are only considering ASW(1) cluster to estimate the BE of the 6 radicals except for OH, we want to examine the effect on the average BE when considering all 9 ASW clusters. For OH, we have both the average BE considering all 9 ASW clusters and the maximum BE from ASW(1) cluster. Hence, we calculate a ratio of 1.177, indicating around deviation, obtained from OH radical calculation by dividing the BSSE-corrected average BE (4287 K) by the BSSE-corrected highest BE (3643 K) of ASW(1) cluster (see LABEL:tab:BE_OH). Consequently, we apply this scaling factor to the other 6 radicals to determine the effect on the average BE over all the 9 ASW clusters, as we only estimated the BEs for these species using the ASW(1) cluster. As mentioned earlier, the standard deviation ( 566 K) for OH is calculated using the Gaussian distribution over all the 9 highest BEs obtained for the 9 ASW clusters, and thus 566 K can be treated as an error bar for the BE of OH. To get an idea of this error bar for the other 6 radicals, a standard deviation ( 786 K) for OH is calculated by taking the 7 binding sites of the ASW(1) cluster. Finally, a scaling factor of is obtained by taking their ratio. For the other 6 radicals, the standard deviation is obtained similarly as for OH radical considering the 7 binding sites of the ASW(1) cluster and then multiplied with the scaling factor of to get an estimate of the standard deviation considering the 9 ASW clusters (see Fig. 11). It is to be noted that the calculation does not take into account the significantly higher BE values for CH and CN, as these are considered special cases. The final BSSE-corrected scaled BE values, along with the error bars, are noted in the last column of Table 2.
Sl. | Radicals | Formula | Ground state | BE in Kelvin | ||||
@@
1. | Methylidyne radical | CH | Doublet | — | — | |||
2. | Imidogen | NH | Triplet | 1947 | — | |||
3. | Hydroxyl radical | OH | Doublet | 3183 | ||||
4. | Mercapto radical | SH / HS | Doublet | 2221 | — | |||
5. | Cyano radical | CN | Doublet | 1736 | — | |||
6. | Nitrogen monosulfide | NS | Doublet | 2774 | — | |||
7. | Nitric oxide | NO | Doublet | — |
@@ 3.2 BSSE correction
The basis set superposition error (BSSE) is an intuitively highly plausible concept. The BSSE correction using the CPC method accounts for the superposition error caused by the sloppy treatment of the basis set for each monomer of a big cluster, as the intermolecular distance varies. Eventually, even if this inconsistency could be eradicated entirely, errors related to the incompleteness of the basis set, known as ?basis set completeness error? (BSCE), would be preserved. Hypothetically, the BSSE and the BSCE would be zero in the complete basis set (CBS) limit. The BSSE and BSCE often have opposite signs (Dunning, 2000). For small and medium basis sets around the vdW minimum, the (negative) BSCE cancels (part of) the (positive) BSSE (Sheng et al., 2011). Ignoring one basis set fault while accounting for the other can produce imbalanced results. According to Sheng et al. (2011), diffuse basis functions can help reduce the BSSE. However, they also pointed out that diffuse basis functions can reduce the BSCE. Therefore, the partial cancellation of these errors still applies, and the extensive use of diffuse basis functions does not lead to a different conclusion. Jensen (2024) recently pointed out that the most significant source of error for smaller basis sets is BSCE. Furthermore, he noted that these errors are dominant in dispersion-type interactions and that including CPC may, in such cases, lead to a deterioration of the results.
The recent studies by Perrero et al. (2022); Bovolenta et al. (2022); Tinacci et al. (2023) suggest that the BE values can be erroneous for bigger systems without the BSSE correction. Considering both the contradictory statements and also to reduce the computational cost and make the calculations possible, we have used only one small basis set, 6-311+G(d,p), and thus cannot extrapolate toward the CBS limit, hence an estimation has been made to determine the influence of BSSE on the highest physisorption BEs of the radicals.
@@
1. | CH | 1892 | 1737 | 1.0892 |
2. | NH | 2083 | 1898 | 1.0975 |
3. | OH | 3990 | 3643 | 1.0952 |
4. | SH / HS | 2416 | 2109 | 1.1456 |
5. | CN | 1192 | 1086 | 1.0976 |
6. | NS | 2606 | 2225 | 1.1712 |
7. | NO | 931 | 598 | 1.5569 |
In Table 3, we compare the results obtained by considering BSSE correction using the CPC method and BSSE-uncorrected BE for the -H binding sites with the highest BE found for all the 7 radicals with ASW(1) cluster. It is to be noted that for the CH radical, the highest BEs are due to chemisorption process (see Table 7). Chemisorption refers to forming a chemical bond between two atoms of the radical and water cluster. The structural geometries (see Fig. 9) suggest that for CH radical, sites 5 and 7 in ASW(1) cluster, the cluster geometries get broken, and consequently, CH2OH and CHOH+H formation happens respectively. It is pretty similar to the case found for the C and CH in Wakelam et al. (2017); Shimonishi et al. (2018); Molpeceres et al. (2021). We consider the highest BE value among only the physisorbed energy values to keep calculation consistency. The result is less evident for CN radical (exceptionally large BSSE-corrected BE of 6317 K with site 5 in ASW(1) cluster) as no structural change can be seen. It is thought that CN forms a non-classical 2c-3e bond (hemibond) between the unpaired electron on the CN carbon atom and the lone pair electrons of the oxygen atom of the water surface in place of forming H-bonds with the surface (Rimola et al., 2018; Martínez-Bachs et al., 2024) and is most probably the reason behind the significantly large BE value. For the case of CN, we consider 1086 K as the BSSE-corrected highest BE value for consistency. To better understand the dependency of BSSE on the BE, the BSSE-corrected and uncorrected BEs are plotted side by side and compared along with the ratio by which the BSSE-corrected BEs are lower than the uncorrected one for all the 7 radicals in Fig. 3. From Table 3, we find that the average ratio , considering the -H binding sites with the highest BE found for all 7 radicals with ASW(1) cluster. A quantitative study on the BSSE dependency of the OH radical considering all the binding sites in ASW(1) cluster is also performed and tabulated in Table 8. The BSSE-corrected and uncorrected values and uncorrected to corrected ratio are plotted in Fig. 10. The uncorrected to corrected average ratio for OH is . It is seen that the BE values with BSSE adjustments are consistently lower than those without BSSE corrections for all the cases. Unless otherwise stated, all the BEs reported or compared afterward are the BSSE-corrected BEs.
@@ 3.3 Comparison of the BEs with previous studies
desorption (TPD), the radical has an enormous chance to diffuse and react, making the experimental evaluation of the desorption parameters almost impossible.
@@
To the best of our knowledge, no experimental studies are available to provide the BEs of radicals considered in this study, except for a handful of experimental determinations for the BE of OH on amorphous silicate surfaces.
@@
Moreover, comparing the theoretical BEs with the experiment is not straightforward.
We must, therefore, rely on the other theoretical existing reference values.
In Table 2, we compare our calculated BE with other available experimental, computational, and public database values.
@@
To use a coherent set of homogeneous data for comparison, we chose the recommended BE values by Das et al. (2018) except for CH and Penteado et al. (2017) and frequently used the KIDA database for all the considered radicals.
Penteado et al. (2017) presented a list of the recommended BEs, which includes data from previous studies and an uncertainty range for each value.
@@
For all the radicals considered here except OH, the provided average BEs are based on the work of Hasegawa & Herbst (1993); Aikawa et al. (1996).
OH: The binding properties of the OH radical are difficult to study experimentally due to its tendency to diffuse and react, leading to the formation of ASW on the bare surface during conventional TPD heating ramps. Only a couple of existing experimental studies reported the BE for hydroxyl radical on bare amorphous silicates as 4600 K (Dulieu et al., 2013) and K (He & Vidali, 2014). Penteado et al. (2017) later recommended BE = K, based on the distribution determined by He & Vidali (2014), taking the average with uncertainty to cover this range.
However, as far as we know, no experimental study considering ASW surface has been published to date.
@@
The OSU gas-phase database provides 2850 K, assuming half the value of (Garrod & Herbst, 2006).
A systematic empirical approach by Wakelam et al. (2017) provided K considering ASW model.
Das et al. (2018) computed by quantum calculations 3183 K considering ASW model.
Ferrero et al. (2020) determined a range of values K considering the amorphous 2D slab ice model, generated from the structures of Shimonishi et al. (2018).
@@
Duflot et al. (2021) found average best estimate of K between minimum 2936 K and maximum 8924 K considering amorphous ice model.
Enrique-Romero et al. (2022) calculated 2911 K and 5376 K on the and ASW ice model.
@@
Very recently, Hendrix et al. (2023) computed the BEs of OH to monomer, dimer, trimer, and tetramer as 1922 K, 3900 K, 4252 K, and 3674 K, respectively using CCSD(T)/aug-cc-pVTZ level of theory and 1907 K, 3905 K, 4433 K, and 3840 K, respectively using DFT-B97X-V/def2-qzvppd level of theory.
They noticed that the H-bond distances between the O atom of OH and the H atom of shorten notably upon adding a molecule to the cluster.
It is seen that from Fig. 4, our calculated average BE of K is well within the range of existing experimental, theoretical, and recommended values and close to the central value.
CH: As per our knowledge, no computational or experimental BE value exists for methylidyne (also called carbyne) radical to date.
Wakelam et al. (2017) attempted to consider CH in their work but noticed the formation of strong (partly) covalent bonding with a single H2O molecule. However, they did not provide any chemisorption value. We calculate its BE considering 7 -H sites of ASW(1) cluster (see Table 7).
@@
A recent study by Nguyen & Peeters (2021) on the reaction rate study of CH with H2O vapor suggests that in the reaction pathway, the first step is the barrier-less association
leading to a pre-reactive complex (PRC), which has a BE of kcal mol-1 ( K). This PRC (HCδ+–δ-OH2) is formed by
@@
a polar interaction of a lone pair orbital of the O atom and an empty p-orbital of the C atom. Subsequently, the PRC can isomerize to CH2OH by insertion of CH into an O–H bond of H2O.
@@
A similar chemisorption behavior of C-atom with H2O was recognized by previous studies (Wakelam et al., 2017; Shimonishi et al., 2018; Molpeceres et al., 2021).
@@
The large dipole moment (1.46 D; Phelps & Dalby, 1966) could be the reason behind the strong long-range interaction and thus showing reactivity toward neutral molecules.
@@
As seen from Table 2 and Fig. 4, our proposed average BE of K is larger by comparing the available estimated (based on educated guesses) range of K provided in Penteado et al. (2017), as well as 925 K provided by the OSU gas-phase database.
@@
To clarify, our average BE is calculated without taking into account the chemisorbed BE values (see Table 7).
@@
NH: We report the average BE for imidogen or nitrene as K, which is in good agreement with the OSU database value of 2378 K and within the range of values of K proposed by Wakelam et al. (2017).
@@
Our calculated value is well above the recommended value of K (Penteado et al., 2017) and larger than the computed value of 1947 K by Das et al. (2018).
Martínez-Bachs et al. (2020) reported the BEs of 4222 K considering crystalline water-ice surface.
@@
Our value agrees well with the best-estimated average value of K considering the amorphous ice model (Duflot et al., 2021).
@@
Enrique-Romero et al. (2022) have recently calculated the BEs of NH as 1560 K and 3909 K on the and ASW ice models respectively considering a single binding site which is indeed a simplistic assumption given that different surface binding sites are available and, accordingly, a distribution of BEs exist.
@@
Enrique-Romero et al. (2022) also noted (from the private communication with B. Martinez-Bachs) the range of BEs of K considering ASW ice surfaces, centered at around 2410 K.
@@
However, the very recent calculations by Martínez-Bachs et al. (2024) show a range of BEs on the ASW model varying from a minimum of 1167 K to a maximum of 4692 K, with an average of 2464 K, which also gives a strong agreement with our value.
@@
SH/HS: We report the average BE for mercapto or sulfanyl radical as K, which is within the range of values of K proposed by Wakelam et al. (2017).
@@
However, the other values 1450 K (OSU database), K (Penteado et al., 2017), 2221 K (Das et al., 2018), and K (Perrero et al., 2022) are found somewhat smaller than our computed average value.
@@
CN: Our computed average BE value for the cyano radical ( K) is consistent with the existing values of 1600 K (OSU database), K (Penteado et al., 2017), and 1736 K (Das et al., 2018), except the larger value K proposed by Wakelam et al. (2017).
@@
However, an exceptionally strong BE of 6317 K is found for site 5 in ASW(1) cluster. This value is not included in the average calculation. Eventually, this high BE value coincides with the range of BEs proposed by Martínez-Bachs et al. (2024), with the values ranging from 2605 K to 7792 K and an average of 4856 K. The hemibond nature is most probably the acceptable reason behind this inconsistency of the BE values.
@@
NS: Unlike the case of the CH radical, our calculated average BE for the nitrogen monosulfide radical K is larger by comparing the available estimated range of K (Penteado et al., 2017), as well as 1900 K provided by the OSU gas-phase database.
Das et al. (2018) calculated value of 2774 K is somewhat close to our calculated average value.
Perrero et al. (2022) calculated the average BE of 2026 K between minimum 1234 K and maximum 2955 K considering an amorphous ice model.
@@
NO: Our computed average BE value of K for nitric oxide radical is well within the range K proposed by Penteado et al. (2017), and the calculated value ( K) on different ASW ice model by Das et al. (2018). The KIDA database suggests a higher value of K as proposed by Wakelam et al. (2017).
@@
Very recently, Hendrix et al. (2023) computed the BEs of NO to monomer, dimer, trimer, and tetramer as 367 K, 1147 K, 438 K, and 1354 K, respectively, using CCSD(T)/aug-cc-pVTZ level of theory and 252 K, 1022 K, 649 K, and 1087 K, respectively, using DFT-B97X-V/def2-qzvppd level of theory. Similar to NH, the results of Martínez-Bachs et al. (2024) give an energy range of 291 K to 965 K, with an average of 712 K, thus strongly agreeing with our results. The weak BE of NO on the surface is primarily governed by dispersive interactions. This is due to the almost negligible dipole moment of NO, which limits the contribution of electrostatic interactions to the overall binding.
It is important to note that the small fluctuations observed in the calculated BEs compared to literature values can be attributed to several factors, including the specific computational method employed, the adopted ice model, and the particular binding sites investigated in this study.

### v2

@@ 3.1 Binding energy
First, we calculated the optimization energy of nine ASW clusters (presented in Table A.1, available on Zenodo), along with their optimized structures. The ASW(1) cluster is, on average, K more stable than the other eight ASW clusters. We then calculated the BE of the OH radical at the -H binding sites of nine ASW clusters. All the calculated (+) BSSE BE values are plotted in Fig. 1 and tabulated in Table A.2 (available on Zenodo).
@@
Instead of taking the average BE over all binding sites, we considered the average over the highest -H binding sites in each cluster. This averaging would be more appropriate for estimating the BE at a low temperature (interstellar condition). We obtain average BE values of K and K for the OH radical with and without considering BSSE corrections, respectively. The locations of the BEs depend on the binding sites and are broadly distributed (see Table A.2, available on Zenodo). To verify the distribution of the BSSE-corrected BE values over the highest -H binding sites in each cluster, we organized them in a distribution plot with a bin width of 500 K (see Fig. 2). The distribution is well reproduced by an overlay of the Gaussian probability distribution function:
@@
where is the mean and is the standard deviation. The fitted parameters for the BE value distribution are the mean () K and standard deviation () K. This implies that the Boltzmann distribution (uniform distribution to the highest BE site at K) holds for each local region, which is further supported by the fitting using the Boltzmann model equation (see Fig. A.1, available on Zenodo). On the contrary, the BE distribution considering all 75 sites (BSSE-corrected and uncorrected) does not mimic a Boltzmann distribution (non-Gaussian type), as depicted in Fig. A.2 (available on Zenodo).
@@
To minimize the computational burden, we performed the calculation of the BE for the remaining six radicals (i.e., CH, NH, SH, CN, NS, and NO) by considering only the ASW(1) cluster. This particular ASW cluster is the most stable (lowest energy) of the nine ASW clusters (see Table A.1, available on Zenodo). Then, we selected the highest physisorbed BE values among the seven binding sites available in ASW(1) (see Table A.3, available on Zenodo). It is interesting to note that all the radicals except NO (i.e., CH, NH, OH, SH, CN, and NS) give the highest BEs (considering both physisorption and chemisorption) with the fifth -H position of the ASW(1) cluster. Since we only considered the ASW(1) cluster to estimate the BE of five of these six radicals (excluding OH), we wanted to examine the effect on the average BE when considering all nine ASW clusters. For OH, we have both the average BE of the nine ASW clusters and the maximum BE from the ASW(1) cluster. We calculate a ratio of 1.177, which indicates a deviation; we got this from the OH radical calculation, by dividing the BSSE-corrected average BE (4287 K) by the highest BSSE-corrected BE (3643 K) of the ASW(1) cluster (see Table A.2, available on Zenodo). Consequently, we applied this scaling factor to the other six radicals to determine the effect on the average BE over all nine ASW clusters, as we only estimated the BEs for these species using the ASW(1) cluster. As mentioned earlier, the standard deviation ( 566 K) for OH was calculated using the Gaussian distribution over the nine highest BEs obtained for the nine ASW clusters, and thus 566 K can be treated as an error bar for the BE of OH. To get an idea of this error bar for the other six radicals, a standard deviation ( 786 K) for OH was calculated by taking the seven binding sites of the ASW(1) cluster. Finally, a scaling factor of was obtained by taking their ratio. For the other six radicals, we obtained the standard deviation like we did for OH radical, using the seven binding sites of the ASW(1) cluster and then multiplying them by the scaling factor of to get an estimate of the standard deviation for the nine ASW clusters (see Fig. A.5, available on Zenodo). It should be noted that this calculation does not take the significantly higher BE values for CH and CN into account, as they are considered special cases. The final BSSE-corrected scaled BE values, along with the error bars, are noted in the last column of Table 2.
| Sl. | Radicals | Formula | Ground-state | BE in Kelvin | ||||
@@
| 1. | Methylidyne radical | CH | Doublet | — | — | |||
| 2. | Imidogen | NH | Triplet | 1947 | — | |||
| 3. | Hydroxyl radical | OH | Doublet | 3183 | ||||
| 4. | Mercapto radical | SH / HS | Doublet | 2221 | — | |||
| 5. | Cyano radical | CN | Doublet | 1736 | — | |||
| 6. | Nitrogen monosulfide | NS | Doublet | 2774 | — | |||
| 7. | Nitric oxide | NO | Doublet | — | ||||
@@ 3.2 BSSE correction
The BSSE is intuitively a highly plausible concept. The BSSE correction using the CPC method accounts for the superposition error caused by the sloppy treatment of the basis set for each monomer of a big cluster, as the intermolecular distance varies. Even if this inconsistency could be eradicated entirely, errors related to the incompleteness of the basis set, known as the basis set completeness error (BSCE), would remain. Hypothetically, the BSSE and the BSCE would be zero in the complete basis set limit. The BSSE and BSCE often have opposite signs (Dunning, 2000). For small and medium basis sets around the van der Waals minimum, the (negative) BSCE cancels (part of) the (positive) BSSE (Sheng et al., 2011). Ignoring one basis set fault while accounting for the other can produce imbalanced results. According to Sheng et al. (2011), diffuse basis functions can help reduce the BSSE. However, they also pointed out that diffuse basis functions can reduce the BSCE. Therefore, the partial cancelation of these errors still applies, and the extensive use of diffuse basis functions does not lead to a different conclusion. Jensen (2024) recently pointed out that the most significant source of error for smaller basis sets is the BSCE. Furthermore, he noted that these errors are dominant in dispersion-type interactions and that including CPC can, in such cases, lead to a deterioration of the results.
The recent studies by Perrero et al. (2022), Bovolenta et al. (2022), and Tinacci et al. (2023) suggest that BE values can be erroneous for bigger systems without BSSE correction. Considering these contradictory statements, and to reduce the computational cost and make the calculations possible, we used only one small basis set, 6-311+G(d,p), and thus cannot extrapolate toward the complete basis set limit. Hence, we relied on an estimate to determine the influence of the BSSE on the highest physisorption BEs of the radicals.
@@
| 1. | CH | 1892 | 1737 | 1.0892 |
| 2. | NH | 2083 | 1898 | 1.0975 |
| 3. | OH | 3990 | 3643 | 1.0952 |
| 4. | SH / HS | 2416 | 2109 | 1.1456 |
| 5. | CN | 1192 | 1086 | 1.0976 |
| 6. | NS | 2606 | 2225 | 1.1712 |
| 7. | NO | 931 | 598 | 1.5569 |
In Table 3 we compare the BE results obtained by carrying out BSSE correction using the CPC method and the BSSE-uncorrected BE for the -H binding sites with the highest BE found for all seven radicals with the ASW(1) cluster. It is to be noted that for the CH radical, the highest BEs are due to the chemisorption process (see Table A.3, available on Zenodo). Chemisorption refers to the formation of a chemical bond between two atoms of the radical and water cluster. The structural geometries (see Fig. A.3, available on Zenodo) suggest that for the CH radical, the cluster geometries get broken, and consequently, CH2OH and CHOH+H formation happens at sites 5 and 7 of the ASW(1) cluster, respectively. This is similar to the case found for C and CH in Wakelam et al. (2017), Shimonishi et al. (2018), and Molpeceres et al. (2021). We considered the highest BE value among only the physisorbed energy values for the sake of calculation consistency. The result is less evident for the CN radical (an exceptionally large BSSE-corrected BE of 6317 K at site 5 in the ASW(1) cluster) as no structural change can be seen. It is thought that CN forms a nonclassical 2c-3e bond (a hemibond) between the unpaired electron on the CN carbon atom and the lone pair of electrons of the oxygen atom of the water surface instead of forming H bonds with the surface (Rimola et al., 2018; Martínez-Bachs et al., 2024) and is most probably the reason for the significantly large BE value. For the case of CN, we considered 1086 K as the highest BSSE-corrected BE value for consistency. To better understand the dependence of the BSSE on the BE, the BSSE-corrected and uncorrected BEs are plotted side by side in Fig. 3 and compared with the ratio by which the BSSE-corrected BEs are lower than the uncorrected one for all seven radicals. From Table 3, we see that the average ratio is , considering the -H binding sites with the highest BE found for all seven radicals with the ASW(1) cluster. A quantitative study on the BSSE dependence of the OH radical taking all the binding sites in the ASW(1) cluster into consideration was also performed (see Table A.4, available on Zenodo). The BSSE-corrected and uncorrected values and the uncorrected-to-corrected ratio are plotted in Fig. A.4 (available on Zenodo). The average uncorrected-to-corrected ratio for OH is . The BE values with BSSE adjustments are consistently lower than those without BSSE corrections for all cases. Unless otherwise stated, all BEs discussed hereafter are BSSE-corrected BEs.
@@ 3.3 Comparison of the BEs with previous studies
desorption, the radical has an enormous chance to diffuse and react, making the experimental evaluation of the desorption parameters almost impossible.
@@
To the best of our knowledge, no experimental studies have provided the BEs of the radicals considered in this study, except for a handful of experimental determinations for the BE of OH on amorphous silicate surfaces.
@@
Moreover, comparing the theoretical BEs with experimentally determined ones is not straightforward.
We must, therefore, rely on existing theoretical reference values.
In Table 2 we compare our calculated BE with other available experimental, computational, and public database values.
@@
To use a coherent set of homogeneous data for comparison, we chose the BE values recommended by (Das et al., 2018, except for CH) and Penteado et al. (2017) and used the KInetic Database for Astrochemistry (KIDA) for many of the radicals.
Penteado et al. (2017) present a list of the recommended BEs, which includes data from previous studies and an uncertainty range for each value.
@@
For all the radicals considered here except OH, the provided average BEs are based on the work of Hasegawa & Herbst (1993) and Aikawa et al. (1996).
OH: The binding properties of the OH radical are difficult to study experimentally due to its tendency to diffuse and react, leading to the formation of ASW on the bare surface during conventional temperature programmed
desorption heating ramps. Only a couple of existing experimental studies have reported the BE for hydroxyl radical on bare amorphous silicates: 4600 K (Dulieu et al., 2013) and K (He & Vidali, 2014). Penteado et al. (2017) later recommended BE = K, based on the distribution determined by He & Vidali (2014), taking the average with uncertainty to cover this range.
However, as far as we know, no experimental study considering ASW surfaces has been published to date.
@@
The Ohio State University (OSU) gas-phase database gives 2850 K, assuming half the value of (Garrod & Herbst, 2006).
A systematic empirical approach by Wakelam et al. (2017) provided K using the ASW model.
Das et al. (2018), using quantum calculations and the ASW model, find 3183 K.
Ferrero et al. (2020) determined the range K using the amorphous 2D slab ice model, generated from the structures from Shimonishi et al. (2018).
@@
Duflot et al. (2021) find an average best estimate of K between minimum 2936 K and maximum 8924 K for the amorphous ice model.
Enrique-Romero et al. (2022) calculated 2911 K and 5376 K for the and ASW ice models.
@@
Very recently, Hendrix et al. (2023) computed the BEs of OH to the monomer, dimer, trimer, and tetramer as 1922 K, 3900 K, 4252 K, and 3674 K, respectively, using the CCSD(T)/aug-cc-pVTZ level of theory and 1907 K, 3905 K, 4433 K, and 3840 K, respectively, using the DFT-B97X-V/def2-qzvppd level of theory.
They noticed that the H-bond distances between the O atom of OH and the H atom of shorten notably when an molecule is added to the cluster.
As shown in Fig. 4, our calculated average BE of K is well within the range of existing experimental, theoretical, and recommended values, and is close to the central value.
CH: As per our knowledge, no computational or experimental BE value exists for the methylidyne (also called carbyne) radical to date.
Wakelam et al. (2017) attempted to consider CH in their work but noticed the formation of strong (partly) covalent bonding with a single H2O molecule. However, they did not provide any chemisorption value. We calculated its BE considering seven -H sites of the ASW(1) cluster (see Table A.3, available on Zenodo).
@@
A recent study by Nguyen & Peeters (2021) on the reaction rate of CH with H2O vapor suggests that in the reaction pathway, the first step is the barrier-less association
leading to a pre-reactive complex, which has a BE of kcal mol-1 ( K). This pre-reactive complex (HCδ+–δ-OH2) is formed via
@@
a polar interaction between a lone pair orbital of the O atom and an empty p-orbital of the C atom. Subsequently, the pre-reactive complex can isomerize to CH2OH if CH is inserted into an O–H bond of H2O.
@@
A similar chemisorption behavior of a C atom with H2O was recognized by previous studies (Wakelam et al., 2017; Shimonishi et al., 2018; Molpeceres et al., 2021).
@@
The large dipole moment (1.46 D; Phelps & Dalby, 1966) could be the reason for the strong long-range interaction, which would mean it is reactive to neutral molecules.
@@
As seen from Table 2 and Fig. 4, our proposed average BE of K is higher than the estimated range of K (based on educated guesses) provided in Penteado et al. (2017), as well as the 925 K provided by the OSU gas-phase database.
@@
To clarify, our average BE was calculated without taking the chemisorbed BE values into account (see Table A.3, available on Zenodo).
@@
NH: We report an average BE for imidogen, or nitrene, of K, which is in good agreement with the OSU database value of 2378 K and within the range K proposed by Wakelam et al. (2017).
@@
Our calculated value is well above the recommended value of K (Penteado et al., 2017) and higher than the value of 1947 K computed by Das et al. (2018).
Martínez-Bachs et al. (2020) report BEs of 4222 K for a crystalline water-ice surface.
@@
Our value agrees well with the best-estimated average value of K for the amorphous ice model (Duflot et al., 2021).
@@
Enrique-Romero et al. (2022) recently calculated the BEs of NH as 1560 K and 3909 K on the and ASW ice models, respectively, considering a single binding site; this is a simplistic assumption given that different surface binding sites are available and, accordingly, a distribution of BEs exists.
@@
Enrique-Romero et al. (2022) also noted (from the private communication with B. Martinez-Bachs) the BE range K for ASW ice surfaces, centered at around 2410 K.
@@
However, the very recent calculations by Martínez-Bachs et al. (2024) show a range of BEs for the ASW model, varying from a minimum of 1167 K to a maximum of 4692 K, with an average of 2464 K, which is also in close agreement with our value.
@@
SH/HS: We report the average BE for mercapto, or the sulfanyl radical, as K, which is within the range K proposed by Wakelam et al. (2017).
@@
However, the other values – 1450 K (OSU database), K (Penteado et al., 2017), 2221 K (Das et al., 2018), and K (Perrero et al., 2022) – are somewhat lower than our computed average value.
@@
CN: Our computed average BE value for the cyano radical ( K) is consistent with the existing values of 1600 K (OSU database), K (Penteado et al., 2017), and 1736 K (Das et al., 2018), but not with the larger value of K proposed by Wakelam et al. (2017).
@@
However, an exceptionally high BE of 6317 K is found for site 5 in the ASW(1) cluster. This value was not included in the average calculation. This high BE value coincides with the range of BEs proposed by Martínez-Bachs et al. (2024): 2605 K to 7792 K, with an average of 4856 K. The hemibond nature is most probably the reason for this inconsistency.
@@
NS: Unlike the case of the CH radical, our calculated average BE for the nitrogen monosulfide radical, K, is higher than the available estimated range of K (Penteado et al., 2017), as well as the 1900 K provided by the OSU gas-phase database.
The value of 2774 K calculated by Das et al. (2018) is somewhat close to our calculated average value.
Perrero et al. (2022) calculated an average BE of 2026 K, a minimum of 1234 K, and a maximum of 2955 K for an amorphous ice model.
@@
NO: Our computed average BE value of K for the nitric oxide radical is well within the range K proposed by Penteado et al. (2017) as well as the range calculated by Das et al. (2018) for a different ASW ice model ( K). The KIDA database suggests a higher value of K as proposed by Wakelam et al. (2017).
@@
Very recently, Hendrix et al. (2023) computed the BEs of the NO to monomer, dimer, trimer, and tetramer as 367 K, 1147 K, 438 K, and 1354 K, respectively, using the CCSD(T)/aug-cc-pVTZ level of theory and 252 K, 1022 K, 649 K, and 1087 K using the DFT-B97X-V/def2-qzvppd level of theory. Similar to NH, the results of Martínez-Bachs et al. (2024) give an energy range of 291 K to 965 K, with an average of 712 K, which strongly agrees with our results. The low BE of NO on the surface is primarily governed by dispersive interactions. This is due to the almost negligible dipole moment of NO, which limits the contribution of electrostatic interactions to the overall binding.